OpenStudy (anonymous):

Part 1: Solve the inequality -3x + 4 less than or equal to x + 8 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1

OpenStudy (anonymous):

-3x + 4 <= x + 8 -3x - x <= 8 - 4 -2x <= 4 (dividing by a negative changes the inequality sign) x >= -2 closed circle on -2, shading to the right.

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

No, thank you

OpenStudy (anonymous):

YW :)